OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4. Could someone please help?

OpenStudy (wintersuntime):

Write an equation for the parabola with focus (4,0) and directrix y=2. Thanks! .. Standard form of parabola: (x-h)^2=4p(y-k), with (h,k) being the (x,y) coordinates of the vertex Given parabola opens downward with axis of symmetry at x=4. Vertex at (4,1). p=1 Equation of parabola: (x-4)^2=-4(y-1) (ans) see the graph below as a visual check on the answer .. y=1-(x-4)^2/4

OpenStudy (anonymous):

I'm so confused I'm sorry @wintersuntime

OpenStudy (wintersuntime):

Are you trying to graph it ?

OpenStudy (wintersuntime):

And it's okay I love helping others

OpenStudy (anonymous):

No. These are the answer I am given: A. y^2 = -8x B. 16y = x^2 C. y = -1/16x^2 D. x = -1/16y^2

OpenStudy (wintersuntime):

What do you think the answer will be ?

OpenStudy (anonymous):

I am thinking C. @wintersuntime

OpenStudy (wintersuntime):

why ?

OpenStudy (anonymous):

Honestly, I'm not sure. I just thought that standard form looked most similar to that choice. @wintersuntime

OpenStudy (anonymous):

its not C

OpenStudy (anonymous):

Could you explain how to find the correct answer then? I am so lost. @saseal

OpenStudy (anonymous):